PLANNING FOR INITIAL BEAM ACTIVITIES IN THE MTA
PLANNING FOR INITIAL BEAM ACTIVITIES IN THE MTA Yagmur Torun, IIT Katsuya Yonehara, Fermilab Dec 16, 2010 – MAP PMG Meeting 1
MTA Status: Detectors & DAQ • Scintillation counters in the hall being retested this week and next • Preparing for DAQ infrastructure to support 2 nd RF station • extra electronics rack installed in Linac Gallery • new patch panels in hall and Linac • new signal cables to be terminated • some parts in the machine shop • Part of the DAQ (CAMAC) that reads out scintillators being debugged • The other part (Lab. View) that controls the RF is stable • Both will be tested with a common external trigger next week during pillbox cavity commissioning • Expect all this to be ready by the end of December • Also want to install some sensors for monitoring the environment in the hall (temp/humidity, air flow and quality, water on floor) • parts mostly in hand, will proceed when other jobs complete 2
MTA Status: Shielding, Cryo, Beamline • Bead installation in penetrations mostly complete • Hand-stack blocks for refrigerator room delivered • Cryo system stable but we have some work to be scheduled • need to replace 2 leaking valves • 2 weeks of downtime required for reconfiguration (to address condensation in the hall) • preferable to do both at the same time • 2 nd compressor hooked up, still to be powered (coordination with Booster) and commissioned • Water from the hall to be dumped outside or moved to MI 65 for evaporation/disposal as necessary • Expect to have cryo manpower by mid-January • New core for 2 nd beam stop is ready, 3 days of work left in the hall to replace the core and install the multiwire • Beam pipe extension to be complete late January • Stability test for beamline magnets carried out last night 3
MTA Status: RF • 805 pillbox is installed and under vacuum • Calibration tomorrow after which we'll be ready to power it • Will also get the surveyors to align it • Instrumentation should be complete next week • Analysis of magnetic field data for the solenoid (needed for box cavity data analysis) in progress • 2 nd 805 MHz RF station set up • RF switch installed • lift table assembled in the hall • waveguide parts available • directional couplers on order • can install HPRF cavity this month 4
MTA Status: Urgent items & issues • We could use help with the following minor issues • Machine shop priority for electronics rack door modifications • Same for beam pipe extension • Tech effort for connector installation on cables in the MTA Hall • Recent retirement program will affect operations • Wickenberg (MTA building manager) • Mau (AD/Operations) • Misek (AD/Mechanical) • Longer term • Preparing for coupling coil installation (early 2012? ) • major task that will require resources 5
Prospect of High Pressurized H 2 gas filled RF cavity project Short term (~ FY 12) • Demonstrate thecavity under high radiation condition • Investigate beam induced plasma in the cavity • Figure out beam loadingmechanismand suppress it • Make a new cavity to test the modification technology • Provide information to high level management group for RF technology selection (including with safety issue) Long term (FY 13 ~) If the HPRF cavity is chosen as the RF technology • Make a realistic RF cavity - Test in cryogenic condition - Test in cooling channel fields - Test gas regulation system - etc 6
Time duration: 3~4 months Nov, 2010 Dec, 2010 Jan, 2011 RF modification, assemble & calibration Time table of HPRF beam test Beam line Wave Guide & Circulator Detector & DAQ Timing Calibration RF test without beam RF beam test 805 MHz button cavity test Beam will be available in January
MTA beam line Final beam absorber Beam profile: MTA experimental hall • Deliver 400 Me. VH- beam in the MTA exp. hall • 1012 to 1013 H-/pulse • Tune beam intensity by using collimator and triplet • Reduce factor from full linac int. down to 1/40 or less MTA solenoid magnet 400 Me. V H- beam MTA exp hall 400 Me. V H- beam 8
Apparatus for beam test • No energized device within 15 feet from HP cavity due to hydrogen safety issue • Beam must be stopped in the magnet due to radiation safety Elastic scattered proton from vacuum window (100~1000 events/pulse) 400 Me. V H- beam New apparatus: • New HPRF cavity • Beam extension line • Collimator+Beam absorber • Luminescence screen+CCD • Beam counter • RF circulator + damper • etc… (Ti) Bound electrons of negative Hydrogen will be fully stripped in the vacuum window Ex) Thickness = 4. 5 g/cm 3 × 0. 1 = 0. 45 g/cm 2 Stripping cross section ~ 10 -18 cm 2 0. 45/47. 9 × 6. 0 1023 = 5. 6 1021 10 -18 = 5600 9
New RF cavity& collimator RF coax line HPRF cell Collimator RF inlet beam Luminescence screen Gas inlet beam Support rail Current status • New RF test cellis ready • New rail systemis ready • New collimatoris ready • Final check and modification 10 10
Beam extension line Elastic scattered proton from vacuum window 400 Me. V H- beam 1 meter 5 meter (Ti) Current status • Discussed with engineer • Final drawing will come out soon • They expected two weeksforconstruction period 11
New beam diagnosticsystem Elastic scattered proton from vacuum window 400 Me. V H- beam (Ti) Current status • Luminescence screen+ CCDare ready • Prepare a telescope beam counter tomake a beam trigger signal • Make a small toroidcoilto measure beam intensity 12
Data flow and acquisition 13
Data acquisition system Name Diagnostic signal ADC input Si. PM/PMT 1 Top plate Light trig. Fast Osc 1 Si. PM/PMT 2 Top plate Spectroscopic Fast Osc 1 3 Si. PM/PMT 3 Topl plate TPB/spare 4 Si. PM/PMT 4 Side wall Light trig. 5 Si. PM/PMT 5 Side wall Spectroscopic 6 Si. PM/PMT 6 Side wall TPB/spare RF Pickup 1 Electric signal Fast Osc 1 RF Pickup 2 Magnetic signal Slow Osc RF Forward 1 Upstream of Circulator RF Return 1 Upstream of Circulator RF Forward 2 Between Circulator&cavity RF Return 2 Between Circulator&cavity Slow Osc Toroid 1 In front of cavity Fast Osc 1 Toroid 2 MTA beamline Fast Osc 2 Beam counter Telescope Fast Osc 2 RF Klystron TTL From MCR Fast Osc 2 Si. PM/PMT 1 2 RF Pickup 1 2 RF Forward 1 RF Return 1 Beam counter System trigger RF Klystron trigger Gate Slow Osc 14 14
Fine timing calibration HPRF Digital Oscilloscope Optical feedthrough ps Laser Si. PM/PMT ½” Heliax for RF pickup signal ½” Heliax for Optical signal Goal: Timing calibration < 100 ps Current status • All devices is ready 15
Summary • Beam test will be ready in January, 2010!! • Study beam induced plasma physics - Made model and simulate RF Q factor reduction due to plasma - Made analysis tool to investigate plasma in the cavity • New RF cavity, collimator system, and beam absorber are ready - Need final check and small modification - Need to build beam extension and vacuum window - Need to extend 800 MHz waveguide - Need to assemble RF power dumper • Diagnostic system - Need to work on beam counter and toroid coil - Luminescence screen and CCD are ready • Data acquisition system -Lab. View system is ready - Need to work on logic signal (CAMAC) system • Documentation for safety assessment - Hydrogen safety document is ready - Preparation of radiation shielding assessment is on going 16
17
18
- Slides: 18